Which of the following are the coordinates of the vertex of y=x^2-10x+2
a. (2,-10)
b. (-5, 23)
c. (5, -23)
d. (-10, 2)

Answer by lwsshak3(11628)

Which of the following are the coordinates of the vertex of
y=x^2-10x+2
complete the square:
y=(x^2-10x+25)-25+2
y=(x-5)^2-23
vertex: (5,-23) ans c.